a) Find all the units of Z[ − 7]. b) Check whether or not < 8x + 6x − 9x + 24 > [x] 3 2 Q is a field. c) Construct a field with 125 element


1
Expert's answer
2021-05-07T09:57:56-0400


Solution:-

(a)


If "a\in R" be a unit and a, b "\in" R , such that


Z(-7) = {a+(-7)b} , a,b "\in" Z


let (a+(-7)) "\in" Z(-7) be a unit "\exist" (c+(-7)d) "\in" (-7)


(a+(-7)b)(c+(-7)d)=1.............(1)


"\implies"

(a-7b)(c-7d)=1


"\implies"

a=1 and c can be 1

as b,d "\in" Z so cannot be fractional





(b)



8x + 6x − 9x + 24 > [x] 3 2 Q is a field


[x] 3 2 Q is a field

on solving we get

x3-2


5x-24 < x3-2

"\implies"

x3-5x+22>0


as we can see


"\boxed{X= -3.3 , 1.6+1.9i , 1.6-1.9i}"



(c)

"Solution: ~We~ know ~ that
\\To ~find ~ an ~ irreducible ~ polynomial ~ of ~ degree ~3 ~ in ~Z_5[x].
\\x^3 +x+1~ is ~ one ~ such ~polynomial~, it~ clearly ~has ~ no ~ linear~ factors~(Since ~0,1,2,3,4,) ~are ~\\not ~roots. 
\\So, F=\frac{Z_5[x]}{x^3 +x+1} ~is ~ a~ field ~ with ~ 5^3 ~elements."